A

Preface of the editor

{ Was hier folgt, will kein persönlicher Beitrag sein, denn als letzter der Letzten, sozusagen nicht Existierender, bin ich mir meiner Wenigkeit und Unwissenheit wohl bewußt. Diskurs über die Schwierigkeit, mit den Toten zu verkehren, Giorgio Manganelli.}

After I found that there is no machine readable version of the AL-GOL 68 Revised Report in the WWW available, I volunteerd to invest some work to make it available. It is hoped that this is accepted as a reasonable approximation to the original, which has been published in Acta Informatica 5 (1975), Heft 1–3.

The IMHO most valuable parts of the RR were scanned and recognized relatively easily using the OCR program which came with my scanner, despite the poor quality of the printed originals, whereas the diagrams and the parts which are typeset in slanted or italic letters, e.g., the IO description and the programming examples, consumed much more work. Though I hope that these parts are reproduced in not too poor condition.

For now, the text is set with plain TEX, using some macros collected in rrmac.tex. The original references inside the text are made browsable with a suitable TEX previewer, e.g., xdvi from the teTeX-Distribution. It should be possible that these references can also be used in ghostview for the pstscript versions, or even with acroread when I find out how to produce a .pdf-file. It is intended, however, to make a pure HTML-version after the quality of the text is sufficiently stabilized.

Some of the original font selections are not truly reproduced. Following the design of TEX I preferred to set those parts containing subscripts in TEX’s math mode. Therefore I typesetted in Euler, what is sans-serif in the original. The reference language is shown in cmit, cmb and cmtt fonts, for distinction of TAG-symbols, TAB-symbols and string-items. Following the most popular stropping regime, I used uppercase for bold-TAGs in the source, however. The mapping is done by a virtual font algol.vf, which is therefore needed to process the document with TEX. The disadvantage of virtual fonts or any use of fonts other than CM family is of course, that the text cannot be typesetted at every TEX-installation. I would like to hear opinions on this. Personally, I feel happier with notions in sans-serif-demibold, as cmbx and cmb fonts don’t look well in metanotions. The serifs are too heavy.

The different levels of the Report’s metalanguage are hopefully clearly visible: Plain description, pragmatic remarks, citations, cross-references, metarules, hyper-rules, protonotions in single and double quotes, paranotions, reference language, modified reference language for transput values (˝ this␣is␣a␣string˝ , 7.07_{10} 1, intrinsic values (3.14, true/, empty/) and the special symbols used in PART V {10.1.3} (⟨ …⟩, ℵ, ℓ, L, C etc.) to compensate for ALGOL 68’s lack of polymorphism and encapsulation features.

Significant help came from C.H.Lindsey, who did proofread early versions, and made a lot of very hard to find corrections. He helped further in understand some structures in the text.

Remarks by C.H.L.Lindsey on my first version of the RR are found in

chl.mbox.

The availability of this text lends to some more experiments which may be interesting, or not. The first which comes to my mind is to see the Report stripped off the pragmatic remarks.

Some – possibly heretic – remarks on this may be found in rant.tex.

{ …, sind die Würmer gezählt und das Haarkleid geprüft, was sollte dann wohl noch übrigbleiben, das der Rede wert wäre? Diskurs …, Giorgio Manganelli.}

What’s left? This work is copyrighted {copylefted} by

    (2000) Wilhelm Bernhard Kloke,  http://vestein.arb-phys.uni-
    dortmund.de,                          //mailto:klokew@acm.org.

It may be used for anything otherwise legal, as long as the above copyright is not removed and that of the original authors of the report is not violated.

{ Ich habe fertig.}